OpenStudy (anonymous):

\[c/9^2 \div c^2/9^3\] divide and simplify.

OpenStudy (anonymous):

we change the operation from divide to multiply and flip the second fraction over (i.e. take its inverse) c 9^3 --- * ----- 9^2 c^2 Now cancelling out common factors from numerator and denominator, we get 9 ---- c

OpenStudy (anonymous):

so you flip it and then cancel out ?

jhonyy9 (jhonyy9):

c/9^2/c^2/9^3=c/9^2 * 9^3/c^2 = 9/c

OpenStudy (anonymous):

harkirat it says incorrect?

OpenStudy (anonymous):

you know why it is totally my fault. it is a not 9

OpenStudy (anonymous):

so I changed the answer to a/c and it was correct! I am so sorry.

OpenStudy (anonymous):

so it was a and not 9????

OpenStudy (anonymous):

yea! When I wrote the problem down it I made the stem too long on the a!

OpenStudy (anonymous):

:)) no problem....